Put the equation in the form ax2 + bx + c = 0. Enter exponents using the caret ( ^ ). For example, you would enter 4x2 as 4x^2.
2x^2+3x=(3x-1)^2-x^2

Answer by rothauserc(4718)

2x^2 +3x = (3x-1)^2 -x^2
2x^2 +3x = 9x^2 -6x +1 -x^2
3x^2 +9x = 9x^2 +1
6x^2 -9x +1 = 0